OpenStudy (anonymous):

an equation that defines y as a function f of x is given a.) solve the equation for y in terms of x, and replace y with with the function f(x) b.)find f(1) x+4y=10

OpenStudy (anonymous):

x+4y=10 -> y = -x/4 + 10/4 <=>f(x) = -x/4 + 10/4 => f(1) = -1/4 + 10/4 = 9/4